# Towards General Neural Surrogate Solvers with Specialized Neural Accelerators

### v1

@@ Abstract
Surrogate neural network-based partial differential equation (PDE) solvers have the potential to solve PDEs in an accelerated manner, but they are largely limited to systems featuring fixed domain sizes, geometric layouts, and boundary conditions. We propose Specialized Neural Accelerator-Powered Domain Decomposition Methods (SNAP-DDM), a DDM-based approach to PDE solving in which subdomain problems containing arbitrary boundary conditions and geometric parameters are accurately solved using an ensemble of specialized neural operators. We tailor SNAP-DDM to 2D electromagnetics and fluidic flow problems and show how innovations in network architecture and loss function engineering can produce specialized surrogate subdomain solvers with near unity accuracy. We utilize these solvers with standard DDM algorithms to accurately solve freeform electromagnetics and fluids problems featuring a wide range of domain sizes.
@@ 1 Introduction
In this work, we propose Specialized Neural Accelerator-Powered Domain Decomposition Methods (SNAP-DDM), which is a qualitatively new way to implement Neural Operators for solving large scale PDE problems with arbitrary domain sizes and boundary conditions. Our method circumvents the issues posed above by subdividing global boundary value problems into smaller boundary value subdomain problems that can be tractably solved with Neural Operators, and then to stitch together subdomain solutions in an iterative, self-consistent method using Domain Decomposition Methods (DDMs). DDMs are the basis for solving large PDE problems with parallel computing (Smith, 1997; Dolean et al., 2015), and they can be implemented using various algorithms including the Schwarz, finite-element tearing and interconnecting (Wolfe et al., 2000), optimized Schwarz (Gander et al., 2002), two-level (Farhat et al., 2000), and sweeping preconditioner (Poulson et al., 2013) methods. While DDM methods have been previously explored in the context of PINNs (Jagtap et al., 2020; Jagtap & Karniadakis, 2021), the accurate solving of arbitrary PDE problems using the combination of Neural Operators and DDM has not been previously reported.
@@ 2.1 SNAP-DDM
2.2 Self-Modulating Fourier Neural Operator
For the subdomain models, we modify the original FNO architecture (Li et al., 2020) and introduce the Self-Modulating Fourier Neural Operator (SM-FNO) subdomain surrogate solver (Figure 2). We specifically incorporate two key features, the first of which we term a modulation encoder. Mechanistically, we utilize multiple residual blocks and fully connected layers to compress the input data into a latent modulation tensor, which then modulates each linear transformation R in the neural operator through element-wise multiplication (Figure 2). This concept builds on our observation that in the original FNO architecture, the linear transform weight R in each Fourier layer is fixed and are independent of the network input parameters, limiting the ability of the neural operator to accurately process the highly heterogeneous input data featured in our problem. This modification is inspired by the efficacy of self-attention in transformer architectures (Vaswani et al., 2017) and multiplicative interactions between inputs in PINNs (Wang et al., 2021a). Other works have also attempted to modify the integral kernel in the FNO layers, with smoothened masking functions that mostly applies to binary input shapesLABEL:liu2024domain. With our grayscale input that represents material dielectric data, our learnable modulation offers more flexibility. We show in Section 4 that the modulation method we introduced is crucial to enhance the expressivity of the model.
@@ 2.3 Hybrid data-physics loss function
Detailed expressions of the data and physics loss terms are in Appendix B. $c$ is a constant weight set to 1 for simplicity (we found the model performance is insensitive to its value between $0.1$ and $10$) and $\alpha$ is a dynamic weighting term that balances data loss and physics loss from history loss statistics. Regularization of network training with physical loss serves to explicitly enforce physical relationships between nearest neighbor fields, which enhances the accuracy of magnetic field spatial derivatives calculated using the finite differences method (Chen et al., 2022b). Such accuracy is critical for evaluating electric fields and Robin boundary conditions from inferenced magnetic fields.
@@ 3.1 Data generation
We used an established finite-difference frequency domain (FDFD) solver (Hughes et al., 2019) to generate 1000 full-wave simulations, each containing heterogeneous dielectric distributions (refractive index from 1 to 4) within a $960\times 960$ grids simulation domain. The structural layouts are specified by a pipeline inspired from image processing, with details provided in Appendix A. Random magnetic current sources surrounding the devices are specified as a superposition of sinusoidal functions with random amplitudes and phases. 40-grid-thick uniaxial PML layers are place on the 4 sides (Gedney, 1996). The resulting simulated fields are then cropped into $64\times 64$ grids sections to produce the subdomain training dataset. Using this approach, we generate a total of 1.2M subdomain training data samples (100k for the PML solver, 100k for the source solver, and 1M for the grayscale material solver).
@@ 3.2 Subdomain network benchmark
We benchmark our trained SM-FNO subdomain solver with the UNet (Ronneberger et al., 2015), Swin Transformer (Liu et al., 2021) (details in Appendix E), the classical FNO, and the recently improved version of FNO termed F-FNO (Tran et al., 2021). We also train our SM-FNO solver without physics loss. The networks are trained with both 100K and 1M total data to show their dependency on data scaling, except for the Swin Transformer model, which is only benchmarked on 100k training data (training on 1M data would take 2 months). The 100k and 1M data are split into 90% training data and 10% test data. All models use a batchsize of 64 and are trained for 100 epochs for 100k training data or 50 epochs for 1M training data. The Adam optimizer with an individually fine-tuned learning rate is used with an exponential decay learning rate scheduler that decreases the learning rate by 30$\times$ by the end of training. A padding of 20 grids is applied to all FNOs.
@@ 3.3 Large scale electromagnetics simulations
Large scale electromagnetic simulations comprising high contrast heterogeneous media are notoriously hard to solve using end-to-end neural surrogate solvers. We show in Appendix F that the training of a Fourier neural operator to solve full-scale problems leads to fundamental scaling bottlenecks in dataset size, model size, and memory usage. We also show in Appendix G that PINNs struggle to scale up to large simulation domains comprising high dielectric contrast media, and that the solutions produced from trained PINN models are particularly sensitive to their detailed initialization and training conditions. These results are consistent with recent large scale simulation demonstrations in the literature: one concept based on graph networks featured errors of 28% (Khoram et al., 2022) and another concept based on neural operators featured errors ranging from 12% to 38% (Gu et al., 2022).
@@ 3.4 Time complexity
In this section, we benchmark the time complexity between SNAP-DDM and a conventional FDFD solver for 2D EM problems. Based on the current accuracy level of SNAP-DDM, we choose to benchmark the time it takes to reach an average mean absolute accuracy of 15%, evaluated on 10 random devices for each domain size. For SNAP-DDM, the number of iterations required for convergence strongly depends on the material refractive index, and we therefore benchmark computation time for simulation domains containing either silicon dioxide (n = 1.5) or titanium dioxide (n = 2.48). Square domains with sizes ranging from $600\times 600$ grids to $2100\times 2100$ grids are evaluated. The FDFD benchmark simulations are performed on 10 grayscale dielectric structures for each domain size with refractive indices ranging from those of silicon dioxide to titanium dioxide. SNAP-DDM is run with one NVIDIA RTX A6000 GPU, and the FDFD solver runs on single CPU of model Intel Xeon Gold 6242R. From the time benchmark, we observe that for silicon dioxide, which has a relatively low refractive index, SNAP-DDM has better performance compared to the FDFD solver. However, for titanium dioxide, SNAP-DDM requires significantly more iterations and ultimately takes a much longer time than FDFD.
@@ 3.5 SNAP-DDM steady state fluid flow simulations
We solve steady state flows in an arbitrary-shaped pipe with circularly shaped obstacles and a viscosity of $\nu=0.08$ (Chen & Doolen, 1998). To train our subdomain boundary value solvers for these problems, we first simulate flows using the time domain Lattice-Boltzmann Method and run the simulations until the flows are at steady state. A total of 200 ground truth simulations with $900\times 600$ grids are generated, from which the data for 100k subdomains with $64\times 64$ grids is produced as the subdomain training dataset. Further details are in Appendix A. The subdomain model takes an image of the obstacle and velocity field ($u,v$) Robin boundary conditions as inputs, and it outputs images of the full velocity field. Ground truth pressures are used with the steady state version of Equation (3) to compute physics loss.
@@ 5 Limitations and future work
Appendix A Data generation
@@
Appendix B Loss functions
$L_{bc}=\frac{1}{N}\sum_{n=1}^{N}\Big{|}\Big{|}g-\big{(}jk(\mathbf{r})\mathbf{H}^{(n)}-\frac{\partial\mathbf{H}^{(n)}}{\partial n}\big{)}\Big{|}\Big{|}_{1}$ (6)
in which $\hat{\mathbf{H}}$ is the ground-truth magnetic field, $k(\mathbf{r})=2\pi\varepsilon(\mathbf{r})/\lambda$ is the wave vector in the medium.
Appendix C Hybrid data-physics training scheme
@@
Appendix D Implementation details for U-Net
@@
Appendix E Implementation details for Swin Transformer
@@
Appendix F End-to-end FNO on the full problem
@@
Appendix G End-to-end PINN on the full problem

### v2

@@ Abstract
Surrogate neural network-based partial differential equation (PDE) solvers have the potential to solve PDEs in an accelerated manner, but they are largely limited to systems featuring fixed domain sizes, geometric layouts, and boundary conditions. We propose Specialized Neural Accelerator-Powered Domain Decomposition Methods (SNAP-DDM), a DDM-based approach to PDE solving in which subdomain problems containing arbitrary boundary conditions and geometric parameters are accurately solved using an ensemble of specialized neural operators. We tailor SNAP-DDM to 2D electromagnetics and fluidic flow problems and show how innovations in network architecture and loss function engineering can produce specialized surrogate subdomain solvers with near unity accuracy. We utilize these solvers with standard DDM algorithms to accurately solve freeform electromagnetics and fluids problems featuring a wide range of domain sizes. Code for this project could be found at: https://github.com/ChenkaiMao97/SNAP-DDM
@@ 1 Introduction
In this work, we propose Specialized Neural Accelerator-Powered Domain Decomposition Methods (SNAP-DDM), which is a qualitatively new way to implement Neural Operators for solving large scale PDE problems with arbitrary domain sizes and boundary conditions. Our method circumvents the issues posed above by subdividing global boundary value problems into smaller boundary value subdomain problems that can be tractably solved with Neural Operators, and then to stitch together subdomain solutions in an iterative, self-consistent method using Domain Decomposition Methods (DDMs). DDMs are the basis for solving large PDE problems with parallel computing (Smith, 1997; Dolean et al., 2015), and they can be implemented using various algorithms including the Schwarz, finite-element tearing and interconnecting (Wolfe et al., 2000), optimized Schwarz (Gander et al., 2002), two-level (Farhat et al., 2000), and sweeping preconditioner (Poulson et al., 2013) methods. We leave more discussions and references of DDM in Appendix A for interested readers. While DDM methods have been previously explored in the context of PINNs (Jagtap et al., 2020; Jagtap & Karniadakis, 2021), the accurate solving of arbitrary PDE problems using the combination of Neural Operators and DDM has not been previously reported.
@@ 2.1 SNAP-DDM
2.2 Self-Modulating Fourier Neural Operator
For the subdomain models, we modify the original FNO architecture (Li et al., 2020) and introduce the Self-Modulating Fourier Neural Operator (SM-FNO) subdomain surrogate solver (Figure 2). We specifically incorporate two key features, the first of which we term a modulation encoder. Mechanistically, we utilize multiple residual blocks and fully connected layers to compress the input data into a latent modulation tensor, which then modulates each linear transformation R in the neural operator through element-wise multiplication (Figure 2). This concept builds on our observation that in the original FNO architecture, the linear transform weight R in each Fourier layer is fixed and are independent of the network input parameters, limiting the ability of the neural operator to accurately process the highly heterogeneous input data featured in our problem. This modification is inspired by the efficacy of self-attention in transformer architectures (Vaswani et al., 2017) and multiplicative interactions between inputs in PINNs (Wang et al., 2021a). Other works have also attempted to modify the integral kernel in the FNO layers, with smoothened masking functions that mostly applies to binary input shapes(Liu et al., 2024). With our grayscale input that represents material dielectric data, our learnable modulation offers more flexibility. We show in Section 4 that the modulation method we introduced is crucial to enhance the expressivity of the model.
@@ 2.3 Hybrid data-physics loss function
Detailed expressions of the data and physics loss terms are in Appendix B. $c$ is a constant weight set to 1 for simplicity (we found the model performance is insensitive to its value between $0.1$ and $10$) and $\alpha$ is a dynamic weighting term that balances data loss and physics loss from history loss statistics (details in Appendix D). Regularization of network training with physical loss serves to explicitly enforce physical relationships between nearest neighbor fields, which enhances the accuracy of magnetic field spatial derivatives calculated using the finite differences method (Chen et al., 2022b). Such accuracy is critical for evaluating electric fields and Robin boundary conditions from inferenced magnetic fields.
@@ 3.1 Data generation
We used an established finite-difference frequency domain (FDFD) solver (Hughes et al., 2019) to generate 1000 full-wave simulations, each containing heterogeneous dielectric distributions (refractive index from 1 to 4) within a $960\times 960$ grids simulation domain. The structural layouts are specified by a pipeline inspired from image processing, with details provided in Appendix C. Random magnetic current sources surrounding the devices are specified as a superposition of sinusoidal functions with random amplitudes and phases. 40-grid-thick uniaxial PML layers are place on the 4 sides (Gedney, 1996). The resulting simulated fields are then cropped into $64\times 64$ grids sections to produce the subdomain training dataset. Using this approach, we generate a total of 1.2M subdomain training data samples (100k for the PML solver, 100k for the source solver, and 1M for the grayscale material solver).
@@ 3.2 Subdomain network benchmark
We benchmark our trained SM-FNO subdomain solver with the UNet (Ronneberger et al., 2015)(details in Appendix E), Swin Transformer (Liu et al., 2021) (details in Appendix F), the classical FNO, and the recently improved version of FNO termed F-FNO (Tran et al., 2021). We also train our SM-FNO solver without physics loss. The networks are trained with both 100K and 1M total data to show their dependency on data scaling, except for the Swin Transformer model, which is only benchmarked on 100k training data (training on 1M data would take 2 months). The 100k and 1M data are split into 90% training data and 10% test data. All models use a batchsize of 64 and are trained for 100 epochs for 100k training data or 50 epochs for 1M training data. The Adam optimizer with an individually fine-tuned learning rate is used with an exponential decay learning rate scheduler that decreases the learning rate by 30$\times$ by the end of training. A padding of 20 grids is applied to all FNOs.
@@ 3.3 Large scale electromagnetics simulations
Large scale electromagnetic simulations comprising high contrast heterogeneous media are notoriously hard to solve using end-to-end neural surrogate solvers. We show in Appendix H that the training of a Fourier neural operator to solve full-scale problems leads to fundamental scaling bottlenecks in dataset size, model size, and memory usage. We also show in Appendix I that PINNs struggle to scale up to large simulation domains comprising high dielectric contrast media, and that the solutions produced from trained PINN models are particularly sensitive to their detailed initialization and training conditions. These results are consistent with recent large scale simulation demonstrations in the literature: one concept based on graph networks featured errors of 28% (Khoram et al., 2022) and another concept based on neural operators featured errors ranging from 12% to 38% (Gu et al., 2022).
@@ 3.4 Time complexity
In this section, we benchmark the time complexity between SNAP-DDM and a conventional FDFD solver for 2D EM problems. Based on the current accuracy level of SNAP-DDM, we choose to benchmark the time it takes to reach an average mean absolute accuracy of 15%, evaluated on 10 random devices for each domain size. For SNAP-DDM, the number of iterations required for convergence strongly depends on the material refractive index, and we therefore benchmark computation time for simulation domains containing either silicon dioxide (n = 1.5) or titanium dioxide (n = 2.48). Square domains with sizes ranging from $600\times 600$ grids to $2100\times 2100$ grids are evaluated. The FDFD benchmark simulations are performed on 10 grayscale dielectric structures for each domain size with refractive indices ranging from those of silicon dioxide to titanium dioxide. SNAP-DDM is run with one NVIDIA RTX A6000 GPU, and the FDFD solver runs on single CPU of model Intel Xeon Gold 6242R. From the time benchmark, we observe that for silicon dioxide, which has a relatively low refractive index, SNAP-DDM has better performance compared to the FDFD solver. However, for titanium dioxide, SNAP-DDM requires significantly more iterations and ultimately takes a much longer time than FDFD. Many aspects of SNAP-DDM can be optimized to improve its efficiency, and we show in Appendix G that by doubling the subdomain size from 64 by 64 grids to 128 by 128 grids, the overall DDM algorithm becomes twice as efficient.
@@ 3.5 SNAP-DDM steady state fluid flow simulations
We solve steady state flows in an arbitrary-shaped pipe with circularly shaped obstacles and a viscosity of $\nu=0.08$ (Chen & Doolen, 1998). To train our subdomain boundary value solvers for these problems, we first simulate flows using the time domain Lattice-Boltzmann Method and run the simulations until the flows are at steady state. A total of 200 ground truth simulations with $900\times 600$ grids are generated, from which the data for 100k subdomains with $64\times 64$ grids is produced as the subdomain training dataset. Further details are in Appendix C. The subdomain model takes an image of the obstacle and velocity field ($u,v$) Robin boundary conditions as inputs, and it outputs images of the full velocity field. Ground truth pressures are used with the steady state version of Equation (3) to compute physics loss.
@@ 5 Limitations and future work
Acknowledgement
This work is funded by National Science Foundation under Award Number 2103301, Office of Naval Research under Award Number N00014-20-1-2105, and 3M under Award Number IC2020-2735.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Reference for DDM methods
The Additive Schwarz Method (ASM) is a simple yet powerful domain decomposition method that is naturally suited for the parallel computation of PDE problems posed using both finite difference and finite element methods. The convergence of the ASM has been theoretically proven for positive-definite systems with elliptical PDEs(Dolean et al., 2015), and it can be accelerated using preconditioners as in the Restricted Additive Schwarz (RAS) method (Cai & Sarkis, 1999), which is used in scientific computation software like PETSc (Balay et al., 1997). For indefinite systems like Maxwell’s Equations, the use of Dirichlet boundary conditions leads to divergence, and Optimized Schwarz Methods (OSM) are required that utilize robin type boundary conditions (Gander, 2006) or higher order and non-local boundary conditions (Gander et al., 2002) to improve convergence.
Beyond the Schwarz Method, two-level and multi-grid methods speed up the convergence by exchanging global information using a global coarse space on top of the subdomain solvers. An example involving the Helmholtz equation is the construction of the Dirichlet-to-Neumann coarse space (Conen et al., 2014), which inherently applies to heterogeneous problems. Other popular methods include Neumann-to-Neumann and Finite element tearing and interconnecting (FETI) methods (Dolean et al., 2015; Farhat & Roux, 1991), which construct a reduced interface system and use Lagrange multipliers to ensure a weak continuity of the solution across the interfaces.
While the current study only uses Neural Network surrogate solvers in the optimized Schwarz Method setup, we believe the efficient solving of general subdomain problems is central to accelerating all classes of domain decomposition methods.
Appendix B Loss functions
$L_{bc}=\frac{1}{N}\sum_{n=1}^{N}\Big{|}\Big{|}g(\mathbf{r})-\big{(}jk(\mathbf{r})\mathbf{H}^{(n)}-\frac{\partial\mathbf{H}^{(n)}}{\partial n}\big{)}\Big{|}\Big{|}_{1}$ (6)
in which $\hat{\mathbf{H}}$ is the ground-truth magnetic field, $g(\mathbf{r})$ is the input boundary value, $k(\mathbf{r})=2\pi\varepsilon(\mathbf{r})/\lambda$ is the wave vector in the medium.
Appendix C Data generation
@@
Appendix D Hybrid data-physics training scheme
@@
Appendix E Implementation details for U-Net
@@
Appendix F Implementation details for Swin Transformer
@@
Appendix G Impact of different subdomain sizes
As pointed out in the main text, the runtime optimization of SNAP-DDM-like systems is a broad topic to be explored in future works. Here we present an immediate runtime improvement by increasing the subdomain size from 64 grids to 128 grids (with the same grid resolution).
We train an 128 by 128 subdomain model with the same SM-FNO-V2 architecture, which in this case has 11.3M weights. We crop the same large simulation data to produce a 1M training dataset of 128 by 128 grids subdomains. The final loss on the material model is 2.06%, about twice as much as our 64 by 64 model, which is expected for the increased subdomain domain size.
We then run the DDM benchmarks shown in the table below, in which we choose 4 different problem sizes, and run DDM algorithm with both 64 and 128 sized models. The iteration and run time are recorded when the averaged MAE loss drops below 10% for 10 random problems with TiO2 as material.
Table 4: DDM benchmark with 64 by 64 grids subdomain model
Problem size Subdomains Iterations Total time (s)
604 by 604 100 37 $2.052\pm 0.006$
968 by 968 256 77 $11.623\pm 0.114$
1208 by 1208 400 111 $26.788\pm 0.164$
1444 by 1444 576 153 $53.734\pm 0.153$
Table 5: DDM benchmark with 128 by 128 grids subdomain model
Problem size Subdomains Iterations Total time (s)
604 by 604 25 18 $0.900\pm 0.001$
968 by 968 64 38 $5.252\pm 0.029$
1208 by 1208 100 52 $11.526\pm 0.064$
1472 by 1472 144 76 $24.771\pm 0.065$
The experiment shows that with the same amount of ground truth data generated, when we scale subdomain model size from 64 to 128, with the same architecture and both trained with 1M data, the subdomain model sacrifice in accuracy, but the DDM algorithm gets twice as fast. An interesting optimization problem arises in balancing subdomain size and model size to achieve the best overall efficiency.
Appendix H End-to-end FNO on the full problem
@@
Appendix I End-to-end PINN on the full problem